OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. y = 7x + 8 y = x + 20 A. (3, 29) B. (2, 22) C. (2, 8) D. (7, 8)

OpenStudy (godlovesme):

plug "7x+8" in the y=x+20 formula for y and try to solve for x

OpenStudy (anonymous):

ok

OpenStudy (godlovesme):

\[7x+8=x+20 \]

OpenStudy (anonymous):

x=2

OpenStudy (godlovesme):

then plug 2 for x and solve for y

OpenStudy (anonymous):

We can eliminate A and D

OpenStudy (godlovesme):

yes :)

OpenStudy (anonymous):

Would it be B?

OpenStudy (godlovesme):

yes u r right :)

OpenStudy (anonymous):

Sweet!! :D Thanks!

OpenStudy (godlovesme):

np